Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 18     Entry time: 05 Feb 2023, 12:32
Author: Natalia, Zhong 
Group: SCS 
Subject: morning shift summary 

achievements:
- Viking spectrometer set up in 8th order around 1280eV and calibrated

problems:
- problems with Viking motors. Took couple of hours to solve, Martin came on site
- problems with moving exit slit (Blade_Bottom motor does not move). Not solved by EEE on call, may be mechanical problems. Called run coordinator, XRO on call (Maurizio) is going to inspect the slit in the tunnel today. Operation of SCS relies on exit slit available (now it opened to ca. 10mm)

ELOG V3.1.4-7c3fd00